How much energy is required to run a 35 W light bulb for 4.0 min?
1
Expert's answer
2020-11-16T07:44:38-0500

P = 35 W

t = 4.0 min = "\frac{4}{60}" = 0.066 h

"E = P \times t"

"E = 35 \times 0.066 = 2.33\; Watt\;hour = 2.33 \times 10^{-3} \; kW\;h"

Proportion:

1 kW-h – "3.6 \times 10^6 \;J"

"2.33 \times 10^{-3} \; kW\;h" – x

"x = \frac{(2.33 \times 10^{-3}) \times (3.6 \times 10^6)}{1} = 8399 \;J"

E = 8400 J